A piston has an external pressure of 14.0atm . How much work has been done if the cylinder goes from a volume of 0.180 liters to 0.530 liters. Express your answer in joules
Answer #1
Concepts and reason

The expansion or compression of gas takes place in a reversible process or an irreversible process based on the conditions used. An irreversible process may occur when internal pressure (Pint)\left( {{{\rm{P}}_{{\rm{int}}}}} \right) is greater than external pressure (Pext)\left( {{{\rm{P}}_{{\rm{ext}}}}} \right) because, the system does not attain equilibrium at every stage of the process.

Fundamentals

Pressure-volume work:

• Irreversible process: In this process, Pin>Pext{{\rm{P}}_{{\rm{in}}}}{\rm{ > }}{{\rm{P}}_{{\rm{ext}}}} and the expansion of gas takes place rapidly without maintaining equilibrium with the surroundings. Therefore, the pressure-volume (PV)\left( {{\rm{PV}}} \right) work done against constant external pressure is given below:

W=PextΔV{\rm{W = }} - {{\rm{P}}_{{\rm{ext}}}}\Delta {\rm{V}}

Where,Wisworkdonebygas,Pextexternalpressure,ΔVischangeinthevolume.\begin{array}{l}\\{\rm{Where, 'W' is work done by gas,}}\\\\ & {\rm{'}}{{\rm{P}}_{{\rm{ext}}}}'{\rm{ external pressure,}}\\\\ & {\rm{'}}\Delta {\rm{V' is change in the volume}}{\rm{.}}\\\end{array}

Given that,

Externalpressure(Pext)=14.0atmInitialvolume(V1)=0.180LitersFinalvolume(V2)=0.530Liters\begin{array}{c}\\{\rm{External pressure}}\left( {{{\rm{P}}_{{\rm{ext}}}}} \right){\rm{ = 14}}{\rm{.0atm}}\\\\{\rm{Initial volume}}\left( {{{\rm{V}}_{\rm{1}}}} \right) = 0.180{\rm{Liters}}\\\\{\rm{Final volume }}\left( {{{\rm{V}}_2}} \right) = 0.530{\rm{Liters}}\\\end{array}

Work done ‘W’ is calculated below:

W=PextΔV=(14.0atm)(0.5300.180)Liters=4.9Litatm\begin{array}{c}\\{\rm{W = }} - {{\rm{P}}_{{\rm{ext}}}}\Delta {\rm{V}}\\\\ = - \left( {{\rm{14}}{\rm{.0atm}}} \right)\left( {0.530 - 0.180} \right){\rm{Liters}}\\\\{\rm{ = }} - 4.9{\rm{Lit}} - {\rm{atm}}\\\end{array}

Convert the units for work done from Lit-atm to joules.

1Litatm=101.3JW(joules)4.9Litatm=496.37Joules\begin{array}{l}\\{\bf{1}}{\rm{Lit - atm = 101}}{\rm{.3J}}\\\\{\rm{W}}\left( {{\rm{joules}}} \right)\\\\ - 4.9{\rm{Lit - atm = }} - {\rm{496}}{\rm{.37Joules}}\\\end{array}

Ans:

Work done if the cylinder goes from a volume of 0.180 liters to 0.530 liters is 496.37J.- 496.37{\rm{J}}{\rm{.}}